# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ Abstract
Quantum phenomena offer the possibility of measuring physical quantities with precision beyond classical limits. However, current progress is constrained by scalability, environmental noise, and challenges in practical integration. This highlights the necessity for novel approaches. An emerging paradigm in this direction is critical quantum metrology—which harnesses the enhanced susceptibility and nonclassical correlations naturally occurring near quantum phase transitions as resources for quantum-enhanced precision. This tutorial provides a pedagogical introduction to key concepts and a detailed overview of prominent quantum sensing strategies that exploit critical phenomena in metrology. Through examples of increasing complexity, the reader is guided through various critical quantum sensing protocols applied to different critical systems. Special emphasis is placed on the optimal scaling of estimation precision with respect to fundamental resources. Finally, we discuss how critical quantum metrology extends from idealized models to realistic open-system and dissipative regimes, and outline both the challenges and opportunities for future quantum technologies.
@@ I Introduction
Over the past few decades, researchers have made remarkable progress in pushing the boundaries of precision beyond the standard quantum limit. Platforms such as trapped ions [15, 16, 17], cold atoms [7, 18, 19, 20, 21, 22], and photonic systems [23, 8, 9] have demonstrated quantum-enhanced measurement techniques, showcasing how quantum mechanics can improve precision. However, these advancements are often restricted to small-scale or mesoscopic quantum systems. The leap to macroscopic systems—where quantum effects could be harnessed on a much larger scale—remains one of the most ambitious challenges in the field. Macroscopic systems face significant hurdles, such as decoherence, technical noise, and the difficulty of generating and maintaining large-scale quantum correlations. Even state-of-the-art technologies like atomic clocks [19], which achieve extraordinary precision, primarily rely on unentangled ensembles of atoms and classical statistical averaging to improve accuracy. Overcoming these barriers requires innovative approaches with superior scalability and greater robustness to noise.
One promising strategy is critical metrology [24, 25, 26], an approach that exploits the extreme sensitivity of quantum systems near critical points of phase transitions. At these critical points, the system statical and dynamical properties becomes highly responsive to external perturbations, providing a natural mechanism for amplifying small changes in parameters of interest. The divergent susceptibility of critical systems is already exploited in scientific and technological sensing applications, such as in bubble chambers [27] and transition edge photodetectors [28]. In particular, critical quantum sensing is an approach that exploits the nonclassical properties developed in proximity of critical points, in order to achieve a metrological quantum advantage [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 48, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 48, 64, 65, 66, 36, 67].
@@ II How to use this document
Before delving into the details, let us briefly outline the structure of this tutorial and how it may best be read. The material is organized to gradually build up from the fundamentals of parameter estimation theory to the forefront of experimental realizations of critical metrology. Readers primarily interested in the core ideas may choose to focus on the sections introducing critical models, while those seeking a broader perspective will find the introductory and concluding parts equally valuable. While every effort has been made to minimize redundancy, some degree of overlap between sections is unavoidable. This is deliberate, as the tutorial is structured to accommodate selective reading: certain concepts are reintroduced in different contexts to ensure coherence and accessibility, regardless of the reader’s chosen path through the material.
Section III offers a concise introduction to parameter estimation theory, covering key notions such as the signal-to-noise ratio, Fisher information, quantum Fisher information and their multiparameter extensions. In Section IV, we explore these concepts in depth using the paradigmatic case of Ramsey interferometry. This example illustrates the standard quantum limit, Heisenberg scaling, the Heisenberg limit, and the influence of decoherence and open-system effects. Readers already familiar with parameter estimation may skip these sections, though we suggest consulting Section IV, as it offers a valuable contrast for the discussion of critical metrology that follows.
@@
Section IX explores driven–dissipative systems, examining both their steady-state and dynamical regimes, and discussing how environmental coupling shapes the prospects for critical metrology under realistic conditions. Section X focuses on ultrastrongly coupled light–matter systems, particularly the quantum Rabi model, as a cutting-edge experimental platform where critical metrological ideas can be directly tested. Together, these sections collocate critical metrology within state-of-the-art experimental realizations, highlighting approaches that capture the essence of critical phenomena while accounting for the inevitable influence of the environment—providing the most realistic pathways toward experimental critical metrology.
Finally, Section XI summarizes the main insights of the tutorial, and Section XII outlines remaining open challenges, and points to promising directions for future research in critical metrology.
@@ III A brief primer on parameter estimation theory
There are two main paradigms for parameter estimation: local and global estimation theory [73, 74, 75, 76, 77, 23, 78]. In local estimation theory, it is assumed that some prior procedure has narrowed the possible values of the parameter to a small interval. The goal is then to determine the parameter value within this local neighbourhood with maximum precision. A widely used tool in this context is the (quantum) Fisher information, which quantifies parameter sensitivity—that is, how well the parameter can be estimated—and the (quantum) Cramér-Rao bound, which sets the fundamental precision limit for any given estimation strategy. In contrast, global estimation theory makes minimal assumptions about prior information, aiming instead to identify or localize the parameter across the entire domain of known possibilities. In this tutorial, our primary focus will be on local estimation theory, and we will use the quantum Fisher information as the main benchmark for precision. For a more comprehensive review of local and global estimation theory, we refer the reader to the seminal works [72, 70] and [72, 79, 80, 81, 82, 83, 84, 85, 86, 87, 78] respectively. In the following section, we provide a brief overview of the key concepts and tools of quantum parameter estimation theory.
Figure 1: Schematic illustration of the parameter estimation procedure for inferring a target parameter $\omega$. A well defined quantum state $\hat{\varrho}$ is initially prepared. The state is then allowed to evolve, during which time information regarding the target parameter is encoded on the quantum state $\hat{\varrho}\!\to\!\hat{\varrho}(\omega)$. After the period of evolution, we perform a measurement whose outcome probabilities are parameter-dependent. After repeating this process $m$ times, we can use these outcomes to build a statistical model from which we can directly infer the parameter of interest by constructing an estimator $\tilde{\omega}$.
@@ III.1 General sensing protocol
Quantum metrology generalizes this same idea to the quantum domain. Instead of a macroscopic liquid, we use well-controlled quantum systems whose states are sensitive to the parameter of interest. In this tutorial, such a probe state is described by a density matrix $\hat{\varrho}$, which may be pure or mixed  [^1] Throughout this tutorial, both operators and matrices are denoted with a hat; the meaning should be clear from the context. . The parameters to be estimated are denoted collectively by $\omega$. With these definitions in place, we can now outline a general quantum sensing protocol. Based on these basic principles, a general sensing protocol can then be defined as follows [see Fig. 1 for a schematic representation]:
@@
  • Evolution, during which the probe interacts with the system and the parameters of interest are dynamically encoded, $\hat{\varrho}\to\hat{\varrho}(\omega)$.
@@
  • Repetition of the previous steps a large number of times, $m$, to construct a statistical dataset, from which the target parameter $\omega$ can be inferred using a suitable estimator, denoted by $\tilde{\omega}$ throughout this manuscript.
@@ III.2 Signal-to-noise ratio
$S_{\omega}\equiv\frac{|\partial_{\omega}\langle\hat{\mathcal{O}}\rangle|^{2}}{\text{Var}(\hat{\mathcal{O}})}=\frac{1}{\text{Var}(\tilde{\omega})}\;$, (1)
@@ III.3 Fisher information
Parameter estimation theory provides a rigorous framework for quantifying the performance of estimating an unknown parameter [70, 73, 72, 74, 71, 90, 91, 92, 14, 1, 2]. In this framework, information about the target parameter $\omega$ is extracted by performing measurements on the parameter-dependent probe state $\hat{\varrho}(\omega)$. We consider a set of measurements $\{\hat{\Pi}_{x}\}$, where each $\hat{\Pi}_{x}$ is a positive operator-valued measure (POVM) element associated with outcome $x$, satisfying the completeness relation $\int_{x}\hat{\Pi}_{x}dx=\mathbb{I}$ for continuous outcomes and $\sum_{x}\hat{\Pi}_{x}=\mathbb{I}$ for discrete outcomes. The conditional probability of obtaining outcome $x$, given that the parameter takes the value $\omega$, is then determined by the Born rule:
@@ III.4 Quantum Fisher information
This provides a clear operational interpretation: The optimal measurement is one whose expectation value is highly sensitive to small changes in $\omega$, while simultaneously minimizing statistical fluctuations. In other words, the optimal measurement basis maximizes the SNR. Although this interpretation is intuitive, a rigorous proof of this identity has only recently been established [94].
@@ III.5 Multiparameter estimation
The most general quantum sensing scenario arises when one seeks to estimate a set of $d$ unknown parameters, $\vec{\omega}=(\omega_{1},\omega_{2},\dots,\omega_{d})$, through suitable measurements of a quantum probe [70, 91, 95]. As in the single-parameter case, information about these parameters is imprinted on the probe state $\hat{\varrho}(\vec{\omega})$. For a given set of POVMs $\{\hat{\Pi}_{x}\}$, the uncertainty in estimating $\vec{\omega}$ from $m$ independent measurements is bounded by the multiparameter Cramér-Rao inequality [72, 70, 71, 74]:
@@
$\hat{\mathcal{I}}_{ij}=\frac{1}{2}\text{Tr}\left[\hat{\varrho}(\vec{\omega})\{\hat{L}_{i},\hat{L}_{j}\}\right]\equiv\text{Tr}\left[\hat{L}_{j}\partial_{\omega_{i}}\hat{\varrho}(\vec{\omega})\right]\;,$ (17)
where $\hat{L}_{i}$ is the SLD operator associated with parameter $\omega_{i}$, defined implicitly by the self-adjoint Lyapunov equation Eq. (9). By construction, the QFIM is a real, symmetric, and positive semidefinite matrix. Notice that the diagonal elements of the QFIM reduce to the single-parameter QFIs, $\hat{\mathcal{I}}_{ii}\equiv\mathcal{I}_{\omega_{i}}$, while the off-diagonal elements capture correlations between parameters. Using the spectral decomposition $\hat{\varrho}(\vec{\omega})=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, the elements of the QFIM can be written in the convenient form [91],
@@
again understood as a matrix inequality. Similarly to the single-parameter case, saturation of this bound requires asymptotically large datasets, prior knowledge of the parameter neighbourhood and the use of optimal unbiased estimators.
@@
We remark that while multiparameter estimation may appear as a straightforward extension of its single-parameter counterpart, important subtleties arise. For instance: optimal measurements for different parameters may be incompatible—preventing simultaneous attainment of bounds; the inherent trade-off in precision due to the statistical correlation between unknown parameters; and the issue of QFIM non-invertibility. These additional complexities can play a significant and subtle role in the context of critical metrology, and we outline their mathematical underpinning in the following.
@@ III.6 Incompatibility problem in multiparameter estimation
In principle, the CRB is always attainable in both single and multiparameter scenarios. By contrast, while the QCRB is always attainable in the single parameter setting, this condition is not generally guaranteed in the multiparameter case. The reason originates from the possible incompatibility of the SLD operators $\hat{L}_{i}$ associated with different parameters $\omega_{i}$ [72, 71, 70, 91, 96, 90, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. If the SLDs are incompatible, no common measurement basis can simultaneously extract the maximal information about all parameters. This phenomenon–known as measurement incompatibility—sets a fundamental limitation, preventing the saturation of the ultimate precision bound in Eq. (20). In such cases, unavoidable trade-off’s between the estimation precisions of different parameters emerge. The necessary and sufficient condition for the attainability of the multiparameter QCRB is
@@
For practical purposes, the problem of measurement incompatibility in the multiparameter setting means that saturation of the QCRB is in general not guaranteed. This has important ramifications, implying that QCRB may underestimate the true estimation error. To properly account for such incompatibilities and define more faithful precision bounds, it is often more appropriate to use the Holevo–Cramér–Rao bound (HCRB) as a benchmark for the ultimate precision limit in multiparameter estimation. The HCRB is defined as [72, 110]
@@
where $W$ is a positive weight matrix—specifying the cost assigned to estimation errors of different parameters. The Holevo bound, $C_{H}\left(W\right)$, is obtained from an optimization problem and is explicitly constructed to incorporate the effects of measurement incompatibility. Importantly, the HCRB is always guaranteed to be asymptotically attainable, even in situations where the QCRB is not. Naturally, when the SLDs are compatible, the issue of measurement incompatibility disappears, and the HCRB reduces to the QCRB. Lastly, it is important to emphasize that evaluation of the HCRB is in general quite challenging—while analytical results are possible for the simplest of models, one typically relies on numerical approaches such as semidefinite programming [95, 111, 112, 99, 113, 114]. Only recently an efficient numerical method, valid for finite dimensional systems, has been proposed [99].
A further subtlety concerns the different assumptions underlying single- versus multiparameter estimation. In the single-parameter setting, one seeks to infer a single target parameter $\omega_{i}$, with the implicit assumption that all other relevant parameters are known with absolute certainty. By contrast, the multiparameter paradigm relaxes this assumption: several parameters $\vec{\omega}$ are to be estimated simultaneously (locally). Consequently, even when a common optimal measurement basis exists and the multiparameter bound is saturable, the precision for any individual parameter is generally lower than that in the corresponding single-parameter scenario. This follows from the matrix inversion in Eq. (15). The effect can be seen straightforwardly in the case of two-parameter estimation, where $\vec{\omega}=(\omega_{1},\omega_{2})$. From Eq. (20), the attainable precision with respect to the parameter $\omega_{i}$ is given by
@@
with $i,j\in\{1,2\}$. In general, this represents a reduction in sensitivity compared to the precision attainable when all other parameters are assumed known, in which case the bound simplifies to $\text{Var}(\tilde{\omega}_{i})\geq 1/\mathcal{I}_{ii}$. The loss of precision is governed by correlations between the multiple unknown parameters, as encoded in the off-diagonal elements of the QFIM. Consequently, strategies designed to optimize the single-parameter QFI are not necessarily optimal in the multiparameter setting; instead, effective approaches must explicitly account for these correlations in order to balance precision across all parameters of interest. In the case of critical quantum probes, this effect can become significantly pronounced as the enhanced sensitivity to parameter variations—arising from diverging susceptibilities near criticality also amplifies the trade-off between parameters [115, 66]. As a result, while criticality can boost the overall metrological potential, it may exacerbate the precision loss for individual parameters.

### v2

@@ Abstract
Quantum phenomena offer the possibility of measuring physical quantities with precision beyond classical limits. However, current progress is constrained by scalability, environmental noise, and challenges in practical integration. This highlights the necessity for novel approaches. An emerging paradigm in this direction is critical quantum metrology—which harnesses the enhanced susceptibility and nonclassical correlations naturally occurring near quantum phase transitions as resources for quantum-enhanced precision. This tutorial provides a pedagogical introduction to key concepts and a detailed overview of prominent quantum sensing strategies that exploit critical phenomena in metrology. Through examples of increasing complexity, the reader is guided through various critical quantum sensing protocols applied to different critical systems. Special emphasis is placed on the optimal scaling of estimation precision with respect to fundamental resources. Finally, we discuss how critical quantum metrology extends from idealized models to realistic open-system, dissipative regimes, and strongly correlated fermionic systems, outlining both the challenges and opportunities for future quantum technologies.
@@ I Introduction
Over the past few decades, researchers have made remarkable progress in pushing the boundaries of precision beyond the standard quantum limit. Platforms such as trapped ions [15, 16, 17], cold atoms [7, 18, 19, 20, 21, 22, 23, 24], and photonic systems [25, 8, 9] have demonstrated quantum-enhanced measurement techniques, showcasing how quantum mechanics can improve precision. However, these advancements are often restricted to small-scale or mesoscopic quantum systems. The leap to macroscopic systems—where quantum effects could be harnessed on a much larger scale—remains one of the most ambitious challenges in the field. Macroscopic systems face significant hurdles, such as decoherence, technical noise, and the difficulty of generating and maintaining large-scale quantum correlations. Even state-of-the-art technologies like atomic clocks [19], which achieve extraordinary precision, primarily rely on unentangled ensembles of atoms and classical statistical averaging to improve accuracy. Overcoming these barriers requires innovative approaches with superior scalability and greater robustness to noise.
One promising strategy is critical metrology [26, 27, 28], an approach that exploits the extreme sensitivity of quantum systems near critical points of phase transitions. At these critical points, the system statical and dynamical properties becomes highly responsive to external perturbations, providing a natural mechanism for amplifying small changes in parameters of interest. The divergent susceptibility of critical systems is already exploited in scientific and technological sensing applications, such as in bubble chambers [29] and transition edge photodetectors [30]. In particular, critical quantum sensing is an approach that exploits the nonclassical properties developed in proximity of critical points, in order to achieve a metrological quantum advantage [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 52, 58, 59, 60, 61, 62, 63, 64, 65, 66, 52, 67, 68, 40, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
@@ II How to use this document
Before delving into the details, let us briefly outline the structure of this tutorial and how it may best be read. The material is organized to gradually build up from the fundamentals of parameter estimation theory to the forefront of experimental realizations of critical metrology, see Tab. 1. Readers primarily interested in the core ideas may choose to focus on the sections introducing critical models, while those seeking a broader perspective will find the introductory and concluding parts equally valuable. While every effort has been made to minimize redundancy, some degree of overlap between sections is unavoidable. This is deliberate, as the tutorial is structured to accommodate selective reading: certain concepts are reintroduced in different contexts to ensure coherence and accessibility, regardless of the reader’s chosen path through the material.
Section III offers a concise introduction to parameter estimation theory, covering key notions such as the signal-to-noise ratio, Fisher information, quantum Fisher information, and their multiparameter extensions. In Section IV, we explore these concepts in depth using the paradigmatic case of Ramsey interferometry. This example illustrates the standard quantum limit, Heisenberg scaling, the Heisenberg limit, and the influence of decoherence and open-system effects. Readers already familiar with parameter estimation may skip these sections, though we suggest consulting Section IV, as it offers a valuable contrast for the discussion of critical metrology that follows.
@@
Section IX explores driven–dissipative systems, examining both their steady-state and dynamical regimes, and discussing how environmental coupling shapes the prospects for critical metrology under realistic conditions. Section X focuses on ultrastrongly coupled light–matter systems, particularly the quantum Rabi model, as a cutting-edge experimental platform where critical metrological ideas can be directly tested. Section XI introduces strongly correlated fermionic systems as a contrasting platform for critical quantum sensing, focusing on an exactly solvable point of the 2 channel Kondo model. Together, these sections collocate critical metrology within state-of-the-art experimental realizations, highlighting approaches that capture the essence of critical phenomena while accounting for the inevitable influence of the environment—providing the most realistic pathways toward experimental critical metrology.
Finally, Section XII summarizes the main insights of the tutorial, and Section XIII outlines remaining open challenges, and points to promising directions for future research in critical metrology.
@@ III A brief primer on parameter estimation theory
There are two main paradigms for parameter estimation: local and global estimation theory [84, 85, 86, 87, 88, 25, 89]. In local estimation theory, it is assumed that some prior procedure has narrowed the possible values of the parameter to a small interval. The goal is then to determine the parameter value within this local neighbourhood with maximum precision. A widely used tool in this context is the (quantum) Fisher information, which quantifies parameter sensitivity—that is, how well the parameter can be estimated—and the (quantum) Cramér-Rao bound, which sets the fundamental precision limit for any given estimation strategy. In contrast, global estimation theory makes minimal assumptions about prior information, aiming instead to identify or localize the parameter across the entire domain of known possibilities. In this tutorial, our primary focus will be on local estimation theory, and we will use the quantum Fisher information as the main benchmark for precision. For a more comprehensive review of local and global estimation theory, we refer the reader to the seminal works [83, 81] and [83, 90, 91, 92, 93, 94, 95, 96, 97, 98, 89], respectively. In the following section, we provide a brief overview of the key concepts and tools of quantum parameter estimation theory.
Figure 1: Schematic illustration of a general parameter estimation protocol. A probe state $\hat{\varrho}$ is initially prepared. The state is then allowed to evolve in time, and the information regarding the target parameter is encoded on the quantum state, $\hat{\varrho}\!\to\!\hat{\varrho}(\omega)$. After the system evolves in time, a measurement—whose outcome probabilities depend on the unknown parameter—is performed. By repeating this process $m$ times, the resulting outcomes can be used to construct a statistical model, from which an estimator $\tilde{\omega}$ can be derived to infer the parameter of interest.
@@ III.1 General sensing protocol
Quantum metrology generalizes this same idea to the quantum domain. Instead of a macroscopic liquid, we use well-controlled quantum systems whose states are sensitive to the parameter of interest. In this tutorial, such a probe state is described by a density matrix $\hat{\varrho}$, which may be pure or mixed  [^1] Throughout this tutorial, both operators and matrices are denoted with a hat; the meaning should be clear from the context. . The parameters to be estimated are denoted collectively by $\omega$. With these definitions in place, we can now outline a general quantum sensing protocol. Based on these basic principles, a general sensing protocol can then be defined as follows (see Fig. 1 for a schematic representation):
@@
  • Evolution, during which the probe interacts with the system, and the parameters of interest are dynamically encoded, $\hat{\varrho}\to\hat{\varrho}(\omega)$.
@@
  • Repetition of the previous steps a large number of times $\mathcal{M}$, to construct a statistical dataset. From this dataset, the target parameter $\omega$ can be inferred using a suitable estimator, denoted by $\tilde{\omega}$ throughout this manuscript.
@@ III.2 Signal-to-noise ratio
$S_{\omega}\equiv\frac{\mathcal{M}\,|\partial_{\omega}\langle\hat{\mathcal{O}}\rangle|^{2}}{\text{Var}(\hat{\mathcal{O}})}=\frac{1}{\text{Var}(\tilde{\omega})}\;$, (1)
@@ III.3 Fisher information
Parameter estimation theory provides a rigorous framework for quantifying the performance of estimating an unknown parameter [81, 84, 83, 85, 82, 101, 102, 103, 14, 1, 2]. In this framework, information about the target parameter $\omega$ is extracted by performing measurements on the parameter-dependent probe state $\hat{\varrho}(\omega)$. We consider a set of measurements $\{\hat{\Pi}_{x}\}$, where each $\hat{\Pi}_{x}$ is a positive operator-valued measure (POVM) element associated with outcome $x$. These elements satisfy the completeness relation $\int_{x}\hat{\Pi}_{x}dx=\mathbb{I}$ for continuous outcomes, and $\sum_{x}\hat{\Pi}_{x}=\mathbb{I}$ for discrete outcomes, where $\mathbb{I}$ denotes the identity operator. The conditional probability of obtaining outcome $x$, given that the parameter takes the value $\omega$, is then determined by the Born rule:
@@ III.4 Quantum Fisher information
This provides a clear operational interpretation: The optimal measurement is one whose expectation value is highly sensitive to small changes in $\omega$, while simultaneously minimizing statistical fluctuations. In other words, the optimal measurement basis maximizes the SNR. Rigorous proofs for this intuitive identity can be found in [105, 106].
@@ III.5 Multiparameter estimation
The most general quantum sensing scenario arises when one seeks to estimate a set of $d$ unknown parameters, $\vec{\omega}=(\omega_{1},\omega_{2},\dots,\omega_{d})$, through suitable measurements of a quantum probe [81, 102, 107]. As in the single-parameter case, information about these parameters is imprinted on the probe state $\hat{\varrho}(\vec{\omega})$. For a given set of POVMs $\{\hat{\Pi}_{x}\}$, the uncertainty in estimating $\vec{\omega}$ from $\mathcal{M}$ independent measurements is bounded by the multiparameter Cramér-Rao inequality [83, 81, 82, 85]:
@@
$\mathcal{I}_{ij}=\frac{1}{2}\text{Tr}\left[\hat{\varrho}(\vec{\omega})\{\hat{L}_{i},\hat{L}_{j}\}\right]\equiv\text{Tr}\left[\hat{L}_{j}\partial_{\omega_{i}}\hat{\varrho}(\vec{\omega})\right]\;,$ (17)
where $\hat{L}_{i}$ is the SLD operator associated with parameter $\omega_{i}$, defined implicitly by the self-adjoint Lyapunov equation Eq. (9). By construction, the QFIM is a real, symmetric, and positive semidefinite matrix. Notice that the diagonal elements of the QFIM reduce to the single-parameter QFIs, $\mathcal{I}_{ii}\equiv\mathcal{I}_{\omega_{i}}$, while the off-diagonal elements capture correlations between parameters. Using the spectral decomposition $\hat{\varrho}(\vec{\omega})=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, the elements of the QFIM can be written in the convenient form [102],
@@
again understood as a matrix inequality. Similarly to the single-parameter case, saturation of this bound requires asymptotically large datasets, prior knowledge of the parameter neighbourhood, and the use of optimal unbiased estimators.
@@
We remark that, while multiparameter estimation may appear as a straightforward extension of its single-parameter counterpart, important subtleties arise. For instance, the optimal measurements for different parameters may be incompatible, preventing the simultaneous saturation of precision bounds; there can be intrinsic trade-offs in precision due to statistical correlations between parameters; and, in some cases, the QFIM might be non-invertible. These additional complexities can play a significant and subtle role in the context of critical metrology, and we outline their mathematical underpinning in the following.
@@ III.6 Incompatibility problem in multiparameter estimation
In principle, the CRB is always attainable in both single and multiparameter scenarios. By contrast, while the QCRB is always attainable in the single parameter setting, this condition is not generally guaranteed in the multiparameter case. The reason originates from the possible incompatibility of the SLD operators $\hat{L}_{i}$ associated with different parameters $\omega_{i}$ [83, 82, 81, 102, 108, 101, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119]. If the SLDs are incompatible, no common measurement basis can simultaneously extract the maximal information about all parameters. This phenomenon, known as measurement incompatibility, sets a fundamental limitation, preventing the saturation of the ultimate precision bound in Eq. (20). In such cases, unavoidable trade-offs between the estimation precisions of different parameters emerge. The necessary and sufficient condition for the attainability of the multiparameter QCRB is
@@
For practical purposes, the problem of measurement incompatibility in the multiparameter setting means that saturation of the QCRB is in general not guaranteed. This has important implications, as the QCRB may then underestimate the true estimation error. To properly account for such incompatibilities and define more faithful precision bounds, it is often more appropriate to use the Holevo–Cramér–Rao bound (HCRB) as a benchmark for the ultimate precision limit in multiparameter estimation. The HCRB is defined as [83, 122]
@@
where $W$ is a positive weight matrix—specifying the cost assigned to estimation errors of different parameters. The Holevo bound, $C_{H}\left(W\right)$, is obtained from an optimization problem and is explicitly constructed to incorporate the effects of measurement incompatibility. Importantly, the HCRB is always guaranteed to be asymptotically attainable, even in situations where the QCRB is not. Naturally, when the SLDs are compatible, the issue of measurement incompatibility disappears, and the HCRB reduces to the QCRB. Lastly, it is important to emphasize that the evaluation of the HCRB is in general quite challenging. While analytical results are possible for the simplest of models, in general one typically relies on numerical approaches, such as semidefinite programming [107, 123, 124, 111, 125, 126]. Only recently an efficient numerical method, valid for finite dimensional systems, has been proposed [111].
A further subtlety concerns the different assumptions underlying single versus multiparameter estimation. In the single-parameter setting, one seeks to infer a single target parameter $\omega_{i}$, with the implicit assumption that all other relevant parameters are known with absolute certainty. By contrast, the multiparameter paradigm relaxes this assumption: several parameters $\vec{\omega}$ are to be estimated simultaneously (locally). Consequently, even when a common optimal measurement basis exists and the multiparameter bound is saturable, the precision for any individual parameter is generally lower than that in the corresponding single-parameter scenario. This follows from the matrix inversion in Eq. (15). The effect can be seen straightforwardly in the case of two-parameter estimation, where $\vec{\omega}=(\omega_{1},\omega_{2})$. From Eq. (20), the attainable precision with respect to the parameter $\omega_{i}$ is given by
@@
with $i,j\in\{1,2\}$. In general, this represents a reduction in sensitivity compared to the precision attainable when all other parameters are assumed known, in which case the bound simplifies to $\text{Var}(\tilde{\omega}_{i})\geq 1/\mathcal{I}_{ii}$. The loss of precision is governed by correlations between the multiple unknown parameters, as encoded in the off-diagonal elements of the QFIM. Consequently, strategies designed to optimize the single-parameter QFI are not necessarily optimal in the multiparameter setting. Instead, effective approaches must explicitly account for these correlations in order to balance precision across all parameters of interest. In the case of critical quantum probes, this effect can become significantly pronounced as the enhanced sensitivity to parameter variations, arising from diverging susceptibilities near criticality, also amplifies the trade-off between parameters [127, 32]. As a result, while criticality can boost the overall metrological potential, it may exacerbate the precision loss for individual parameters.